# Tensor invariants of finite and classical groups

### v1

@@ 4. Some results for invariants of classical groups
4.1. Universal good filtrations
@@
We fix a Borel subgroup $B\subset G$ containing a maximal split torus $T$ of $G$. For a dominant integral weight $\lambda$, there exists a $G$-representation $\nabla(\lambda)$ such that the dual Weyl module of $G_{k}$ corresponding to $\lambda$ is isomorphic to the dual Weyl module $\nabla(\lambda)_{k}$ for every algebraically closed field $k$. We set $\nabla(\lambda)$ to be the dual Weyl module of $G$ corresponding to $\lambda$ over $\mathbf{Z}$. A $G$-representation has a universal good filtration if it has a filtration such that the successive quotients are isomorphic to dual Weyl modules $\nabla(\lambda)$ with $\lambda$ allowed to vary. Every $G$-representation which admits a universal good filtration is free as a $\mathbf{Z}$-module since the dual Weyl modules over $\mathbf{Z}$ are themselves free $\mathbf{Z}$-modules.
@@ 4.2. Results
Assume $W$ is a finite $\mathbf{Z}$-module with an algebraic action of $G$ admitting a universal good filtration. For $p\gg 0$, the invariant ring $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$.
@@
By [DM21, Theorem 1.10(2)], the $\mathbf{GL}$-algebra $\operatorname{Sym}(W_{k}\otimes\mathbf{V})^{G_{k}}$ is finitely generated for $p\gg 0$. So $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$ by Corollary 3.3. ∎
@@
Consider the rank $n^{2}$ free $\mathbf{Z}$-module $W$ consisting of $n\times n$ square matrices endowed with the conjugation action of $\mathbf{GL}_{n}$. The representation $W$ has a good filtration over $\mathbf{Z}$ [DKZ02, Page 63]. However, the tca $T(W_{k})^{GL_{n}(k)}$ is not finitely generated if char $k\leq n$. Indeed, in Corollary 3.2 of loc. cit., the authors show that a minimal generating set of $\operatorname{Sym}(W\otimes\mathbf{V})^{GL_{n}(k)}$ consists of certain trace elements of weight $1^{m}$ for all $m>0$. The result now follows from Lemma 3.4 and Lemma 2.3(2). ∎
@@
Assume $W$ is a $G$-representation over $\mathbf{Z}$ with a universal good filtration. For $p>\dim(W)$, the invariant ring $T(W_{k})^{G_{k}}$ has a Specht filtration.
@@
By [DM21, Corollary 6.8], when $\operatorname{char}(k)>\dim(W)$, the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})^{G_{k}}$ has a filtration with associated graded pieces being the Schur modules $\mathbf{S}_{\lambda}(\mathbf{V})$. The result now follows by combining (1) and (3) of Lemma 2.3. ∎
@@ 5. Gelfand–Kirillov dimension
Recall that $k$ is an infinite field of characteristic $p\geq 0$. Let $\mathcal{C}$ be a symmetric monoidal category over $k$ of Krull–Gabriel dimension zero, and $A$ be a finitely generated algebra object in $\mathcal{C}$ with $V\subset A$ being a finite length subobject of $A$ that generates it. Let $f_{A;V}(n)$ be the length of the image $A_{V,\leq n}$ of the canonical map $\bigoplus_{i\leq n}\operatorname{Sym}^{i}(V)\to A$.
@@
It suffices to show that $\dim(A;V)\leq\dim(A;W)$. There exists $l>0$ such that $W\subset A_{V,\leq l}$, so we have $f_{A;W}(n)\leq f_{A;V}(nl)$ which in turn gives the required inequality. ∎
@@
Recall that $\mathbf{S}_{\lambda}(W)$ vanishes if $\lambda$ has $>\dim(W)$ nonzero parts. For the sake of simplicity, we now assume $\operatorname{char}(k)=p=0$ so that $\operatorname{len}(S^{\lambda})=1$ in the above sum. We now estimate $f_{A;W}$: the number of partitions of $N$ with at most $n$ parts is $\approx cN^{n-1}$ with $c$ a constant for $N\gg n$; and the dimension of the Schur module $S_{\lambda}(W)$ with $|\lambda|\gg n$ and $\lambda_{n+1}=0$ is $\approx d|\lambda|^{\frac{n(n-1)}{2}}$ where $d$ is another constant, as can be seen using Weyl’s dimension formula. Therefore, $f_{A;W}(N)\approx\sum_{i=0}^{N}cdi^{n-1}i^{\frac{n(n-1)}{2}}$ which has dominant term $N^{\frac{n(n+1)}{2}}$, from which we get $\operatorname{GKdim}(T(W))=\binom{n+1}{2}$. ∎
It would be interesting to also compute $\operatorname{GKdim}$ of $T(W)$ when $p>0$. If $\dim(W)\in\{1,2\}$, then $\operatorname{GKdim}$ is independent of characteristic as $\operatorname{len}(S^{\lambda})$ for partitions with $\leq 2$ parts is at most $\approx\log(|\lambda|)$ (see Example 5.6).
@@
On the $\mathbf{GL}$-algebra side, the $\mathbf{SL}_{2}(k)$-invariant subring of $\operatorname{Sym}(\mathbf{V}\oplus\mathbf{V})$ is isomorphic to the quotient of $\operatorname{Sym}({\textstyle\bigwedge}^{2}(\mathbf{V}))$ by the ideal generated by the Plücker relations. By [Bof91, Theorem 2.7], this quotient ring has a multiplicity-free filtration by the Schur modules $\mathbf{S}_{(2^{n})}(\mathbf{V})$. Applying the Schur functor and using Lemmas 3.2 and 2.3(3), we see that $A\coloneqq T(k^{2})^{\mathbf{SL}_{2}(k)}$ has a multiplicity-free filtration by the Specht modules $S_{(2^{n})}$. If $\operatorname{char}(k)=p=0$, we immediately see that $\operatorname{GKdim}(A)=1$. If instead $p>0$, then we have $S^{\lambda}\cong S^{\lambda^{t}}\otimes\operatorname{sgn}$ where $\operatorname{sgn}$ is the sign representation of $\mathfrak{S}_{|\lambda|}$, so $\operatorname{len}(S_{(2^{n})})=\operatorname{len}(S_{(n,n)})$. By the work of James (see [Fay03, Theorem 2.5]), the length of $S_{(n,n)}$ is $\approx\lfloor\log_{p}(\frac{n+1}{2})\rfloor$. Combining the previous two facts, we get that $f_{A;A_{2}}(n+1)-f_{A;A_{2}}(n)\approx\lfloor\log_{p}(\frac{n+1}{2})\rfloor$ from which we also get $\operatorname{GKdim}(A)=1$. ∎
Acknowledgements.
The author thanks Mátyás Domokos, Martin Kassabov and Andrew Snowden for helpful discussions over email, and is especially grateful to Lucas Buzaglo for patiently explaining foundational results about noncommutative rings.

### v2

@@ 4. Some results for invariants of classical groups
4.1. Universal good filtration
@@
We fix a Borel subgroup $B\subset G$ containing a maximal split torus $T$ of $G$. For a dominant integral weight $\lambda$, there exists a $G$-representation $\nabla(\lambda)$ such that the dual Weyl module of $G_{k}$ corresponding to $\lambda$ is isomorphic to the dual Weyl module $\nabla(\lambda)_{k}$ for every algebraically closed field $k$. We set $\nabla(\lambda)$ to be the dual Weyl module of $G$ corresponding to $\lambda$ over $\mathbf{Z}$. A $G$-representation has a universal good filtration if it has a filtration such that the successive quotients are isomorphic to dual Weyl modules $\nabla(\lambda)$ with $\lambda$ allowed to vary. Every $G$-representation that admits a universal good filtration is a free $\mathbf{Z}$-module since the dual Weyl modules over $\mathbf{Z}$ are themselves free $\mathbf{Z}$-modules.
@@ 4.2. Results
Assume $W$ is a finite $\mathbf{Z}$-module with an algebraic action of $G$ that admits a universal good filtration. There exists $N>0$ such that the invariant ring $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$ for all algebraically closed fields $k$ of char $p>N$.
@@
By [DM21, Theorem 1.7(2)] (Theorem 1.10(2) in the arXiv version), the $\mathbf{GL}$-algebra $\operatorname{Sym}(W_{k}\otimes\mathbf{V})^{G_{k}}$ is finitely generated for $p\gg 0$, and so $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$ by Corollary 3.2. ∎
@@
Consider the rank $n^{2}$ free $\mathbf{Z}$-module $W$ consisting of $n\times n$ square matrices endowed with the conjugation action of $\mathbf{GL}_{n}$. The representation $W$ has a good filtration over $\mathbf{Z}$ [DKZ02, Page 63]. However, the tca $T(W_{k})^{GL_{n}(k)}$ is not finitely generated if char $k\leq n$. Indeed, in Corollary 3.2 of loc. cit., the authors show that a minimal generating set of $\operatorname{Sym}(W\otimes\mathbf{V})^{GL_{n}(k)}$ contains certain trace elements of weight $1^{m}$ for all $m>0$. The result now follows from Lemma 3.3 and Lemma 2.3(2). ∎
@@
Assume $W$ is a $G$-representation over $\mathbf{Z}$ with a universal good filtration. For $p>\dim(W)$, the invariant ring $T(W_{k})^{G_{k}}$ has a filtration where the successive quotients are Specht modules.
@@
By [DM21, Corollary 27] (Corollary 6.8 in the arXiv version), when $\operatorname{char}(k)>\dim(W)$, the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})^{G_{k}}$ has a filtration with associated graded pieces isomorphic to Schur modules $\mathbf{S}_{\lambda}(\mathbf{V})$. The result now follows by combining (1) and (3) of Lemma 2.3. ∎
@@ 5. Gelfand–Kirillov dimension
Recall that $k$ is an infinite field of characteristic $p\geq 0$. Let $\mathcal{C}$ be a symmetric monoidal $k$-linear category of Krull–Gabriel dimension zero, and $A$ be a finitely generated algebra object in $\mathcal{C}$ with $V\subset A$ being a finite length subobject of $A$ that generates it. Let $f_{A;V}(n)$ be the length of the image $A_{V,\leq n}$ of the canonical map $\bigoplus_{i\leq n}\operatorname{Sym}^{i}(V)\to A$.
@@
It suffices to show that $\dim(A;V)\leq\dim(A;W)$. There exists $l>0$ such that $W\subset A_{V,\leq l}$, so we have $f_{A;W}(n)\leq f_{A;V}(nl)$, which in turn gives the required inequality. ∎
@@
Recall that $\mathbf{S}_{\lambda}(W)$ vanishes if $\lambda$ has $\dim(W)$ nonzero parts. For the sake of simplicity, we now assume $\operatorname{char}(k)=p=0$ so that $\operatorname{len}(S^{\lambda})=1$ in the above sum. We now estimate $f_{A;W}$: the number of partitions of $N$ with at most $n$ parts is $\approx cN^{n-1}$ with $c$ a constant for $N\gg n$; and the dimension of the Schur module $S_{\lambda}(W)$ with $|\lambda|\gg n$ and $\lambda_{n+1}=0$ is $\approx d|\lambda|^{\frac{n(n-1)}{2}}$ where $d$ is another constant, as can be seen using Weyl’s dimension formula. Therefore, $f_{A;W}(N)\approx\sum_{i=0}^{N}cdi^{n-1}i^{\frac{n(n-1)}{2}}$ which has dominant term $N^{\frac{n(n+1)}{2}}$. So $\operatorname{GKdim}(T(W))=\binom{n+1}{2}$. We emphasize that this agrees with the computation of the Krull–Gabriel dimension of the module category of the tca $T(W)$ in [SS19, § 3.5]. ∎
It would be interesting to also compute $\operatorname{GKdim}$ of $T(W)$ when $p>0$. If $\dim(W)\in\{1,2\}$, then $\operatorname{GKdim}$ is independent of characteristic, as $\operatorname{len}(S^{\lambda})$ for partitions with $\leq 2$ parts is at most $\approx\log(|\lambda|)$ (see references cited in Example 5.6).
@@
On the $\mathbf{GL}$-algebra side, the $\mathbf{SL}_{2}(k)$-invariant subring of $\operatorname{Sym}(\mathbf{V}\oplus\mathbf{V})$ is isomorphic to the quotient of $\operatorname{Sym}({\textstyle\bigwedge}^{2}(\mathbf{V}))$ by the ideal generated by the Plücker relations. By [Bof91, Theorem 2.7], this quotient ring has a multiplicity-free filtration by the Schur modules $\mathbf{S}_{(2^{n})}(\mathbf{V})$. Applying the Schur functor and using Theorem 1.2 and Lemma 2.3(3), we see that $A\coloneqq T(k^{2})^{\mathbf{SL}_{2}(k)}$ has a multiplicity-free filtration by the Specht modules $S_{(2^{n})}$. If $\operatorname{char}(k)=p=0$, we immediately see that $\operatorname{GKdim}(A)=1$. If instead $p>0$, then we have $S^{\lambda}\cong S^{\lambda^{t}}\otimes\operatorname{sgn}$ where $\operatorname{sgn}$ is the sign representation of $\mathfrak{S}_{|\lambda|}$, so $\operatorname{len}(S_{(2^{n})})=\operatorname{len}(S_{(n,n)})$. By the work of James (see [Fay03, Theorem 2.5]), the length of $S_{(n,n)}$ is $\approx{\lfloor\log_{p}(\frac{n+1}{2})\rfloor}$. Combining the previous two facts, we get that $f_{A;A_{2}}(n+1)-f_{A;A_{2}}(n)\approx\log_{p}(n)$ from which we obtain $\operatorname{GKdim}(A)=1$. ∎